# ASGO: Adaptive Structured Gradient Optimization

### v2

@@ C.1 Pretraining NanoGPT
Experimental Setup:
As an initial experiment, we compared the performance of ASGO, DASGO, Muon, Shampoo, and Adam-W for training NanoGPT, which consists of 6 Transformer layers, 6 attention heads, and an embedding dimension of 384, on the Shakespeare character-level dataset with a sequence length of 256 tokens. For all algorithms, we trained the model for 20 epochs, where each epoch contained 128 steps, using a batch size 128 and the OneCycle learning rate (lr) schedule.
Before discussing hyperparameter (HP) tuning, we note that recent research [Shi et al., 2023, Morwani et al., 2024, Lin et al., 2024] has suggested treating the inverse order (IO) in Shampoo’s preconditioner as a tunable HP rather than using the standard value of $-1/4$. Notably, Morwani et al. [2024] demonstrated that Shampoo with IO$=-1/2$ provides a superior approximation of the full-matrix AdaGrad optimizer. Consequently, we treated the IO as a Shampoo HP in our pretraining experiments. Additionally, Shampoo’s performance is highly dependent on the initialization of its preconditioner matrices, as this is crucial for accurately approximating the Empirical Fisher Information Matrix [Morwani et al., 2024]. To address this initialization sensitivity, we implemented a preconditioner warmup phase specifically for Shampoo. Following the methodology in Ren and Goldfarb [2021], we dedicated the first epoch exclusively to accumulating statistics for the preconditioner matrices without updating model parameters. This approach yields a more robust estimation of the preconditioner, which significantly improves the stability of Shampoo during subsequent training iterations. Figure 5 depicts the training loss and validation loss for training NanoGPT with Shampoo, DASGO, Muon, ASGO and Adam-W.
@@
To ensure optimal performance for each optimizer, we conducted a comprehensive HP search using random search [Bergstra and Bengio, 2012, Choi et al., 2019] within the following predefined ranges: initial lr: $[10^{-5},10^{-1}]$; $\beta_{1}$: $[0.7,0.99]$; $\beta_{2}$: $[0.7,0.99]$; and selected the warmup factor from the set $[0.1,0.15,0.2,0.25,0.3]$. For the optimizers Shampoo and ASGO, we selected the update frequency $\tau$ from the set $[5,10,15]$ and for Shampoo, we selected its IO as either $-\frac{1}{2}$ or $-\frac{1}{4}$. HPs were selected based on validation loss after 20 epochs of training. For each optimizer, we performed a random search for 16 hours to find the optimal HP combination. These best HP choices for training NanoGPT are given in Table 3.
@@ C.2 Pretraining GPT2
The model configuration consists of 12 Transformer layers, 12 attention heads, and an embedding dimension of 768. The total number of training steps was 1000, with a batch size per GPU of 32. The model was trained using Distributed Data Parallel (DDP) on 4 NVIDIA V100 GPUs, employing gradient accumulation over 8 steps. To ensure a fair comparison across all optimization algorithms, a consistent learning rate schedule was utilized: a linear warmup for the first 200 steps, followed by a cosine annealing decay to a final learning rate of $1\times 10^{-5}$ for the remainder of the training process. Training was conducted on the OpenWebText dataset with a sequence length of 512 tokens.
@@
  • Learning rate (lr): $[1\times 10^{-5},1\times 10^{-4},5\times 10^{-4},1\times 10^{-3},5\times 10^{-3},1\times 10^{-2},5\times 10^{-2},0.1,0.5]$
@@
Optimizer Learning Rate $\beta_{1}$ $\beta_{2}$ Damping $\epsilon$ Update Freq. ($\tau$)
Muon 0.01 0.9 N/A N/A N/A
AdamW 0.005 0.9 0.99 $1\times 10^{-6}$ N/A
DASGO 0.01 0.9 0.99 $1\times 10^{-6}$ N/A
ASGO 0.1 0.9 0.95 $1\times 10^{-6}$ 1
@@ C.3 Finetune GPT2-Large
For the fine-tuning experiments on WikiText-2, our hyperparameter tuning focused exclusively on identifying the optimal learning rate for each optimizer under both the CLM and FIM objectives. We performed a grid search over a predefined set of learning rate candidates: $[1\times 10^{-6},5\times 10^{-5},1\times 10^{-4},5\times 10^{-4},1\times 10^{-3},5\times 10^{-3}]$. The optimal learning rate for each optimizer and fine-tuning objective combination was determined by selecting the learning rate that yielded the lowest validation perplexity on the WikiText-2 validation set after the full 2 epochs of fine-tuning. Table 5 shows the optimal learning rate.
@@
CLM $5\times 10^{-5}$ $5\times 10^{-4}$ $1\times 10^{-3}$ $1\times 10^{-3}$
FIM $1\times 10^{-4}$ $5\times 10^{-4}$ $1\times 10^{-3}$ $1\times 10^{-3}$
@@
if $\text{if }m<n\text{ then }V_{t}=\beta_{2}V_{t-1}+\left(1-\beta_{2}\right)M_{t}M_{t}^{\top}$ (4)
else $\text{else }V_{t}=\beta_{2}V_{t-1}+\left(1-\beta_{2}\right)M_{t}^{\top}M_{t}$
if $\text{if }m<n\text{ then }W_{t+1}=W_{t}-\eta_{t}(V_{t}+\epsilon I)^{-\frac{1}{2}}M_{t}$
else $\text{else }W_{t+1}=W_{t}-\eta_{t}M_{t}(V_{t}+\epsilon I)^{-\frac{1}{2}}$
@@
  • Modified ASGO Update Direction (Coupled Newton, $\epsilon=10^{-8}$): $\Delta W^{\text{ASGO}}_{\text{CN},\epsilon=10^{-8}}=(V_{t}+10^{-8}I)^{-\frac{1}{2}}M_{t}$, where $(V_{t}+10^{-8}I)^{-\frac{1}{2}}$ is computed using the Coupled Newton algorithm for 50 iteration steps. Higham [2008], Shi et al. [2023]
  • Modified ASGO Update Direction (Newton-Schulz, $\epsilon=10^{-8}$): $\Delta W^{\text{ASGO}}_{\text{NS},\epsilon=10^{-8}}=(V_{t}+10^{-8}I)^{-\frac{1}{2}}M_{t}$, where $(V_{t}+10^{-8}I)^{-\frac{1}{2}}$ is computed using the Newton-Schulz algorithm for 50 iteration steps. [^5] The Newton-Schulz algorithm is derived from Algorithm 6.35 on Page 153 in Higham [2008]. We utilized the quintic version with parameters $(a,b,c)=(2,-1.5,0.5)$ as described in Jordan et al. [2024]. Higham [2008], Bernstein and Newhouse [2024b], Jordan et al. [2024]
@@ Appendix E Proof of Nonsmooth Convergence of ASGO
[…]
ASGO: $\quad\mathcal{O}\left(D_{\mathrm{op}}\cdot{\rm tr}\left(\left(\sum_{t=0}^{T-1}G_{t}G_{t}^{\top}\right)^{\frac{1}{2}}\right)\right)\leq\mathcal{O}\left(D_{\mathrm{op}}\cdot\sum_{j=1}^{m}\sqrt{\sum_{i=1}^{n}\sum_{t=0}^{T-1}[G_{t}]_{i,j}^{2}}\right).$

### v3

@@ C.1 Pretraining NanoGPT
As an initial experiment, we compared the performance of ASGO, DASGO, Muon, Shampoo, and AdamW for training NanoGPT, which consists of 6 Transformer layers, 6 attention heads, and an embedding dimension of 384, on the Shakespeare character-level dataset with a sequence length of 256 tokens. For all algorithms, we trained the model for 20 epochs, where each epoch contained 128 steps, using a batch size 128 and the OneCycle learning rate (lr) schedule. In contrast to our experiments on GPT2, where we trained 1D and Embedding layer parameters as done in AdamW, we trained all the parameters independently in ASGO.
Before discussing hyperparameter (HP) tuning, we note that recent research [Shi et al., 2023, Morwani et al., 2024, Lin et al., 2024] has suggested treating the inverse order (IO) in Shampoo’s preconditioner as a tunable HP rather than using the standard value of 1/4-1/4. Notably, Morwani et al. [2024] demonstrated that Shampoo with IO=1/2=-1/2 provides a superior approximation of the full-matrix AdaGrad optimizer. Consequently, we treated the IO as a Shampoo HP in our pretraining experiments. Additionally, Shampoo’s performance is highly dependent on the initialization of its preconditioner matrices, as this is crucial for accurately approximating the Empirical Fisher Information Matrix [Morwani et al., 2024]. To address this initialization sensitivity, we implemented a preconditioner warmup phase specifically for Shampoo. Following the methodology in Ren and Goldfarb [2021], we dedicated the first epoch exclusively to accumulating statistics for the preconditioner matrices without updating model parameters. This approach yields a more robust estimation of the preconditioner, which significantly improves the stability of Shampoo during subsequent training iterations. Figure 4 depicts the training loss and validation loss for training NanoGPT with Shampoo, DASGO, Muon, ASGO and Adam-W.
Examining both test and training loss curves, ASGO consistently outperforms Shampoo despite requiring only half the memory consumption and computational effort, which highlights ASGO’s practical advantages for training language models. Furthermore, ASGO and Muon achieve the lowest final training and test losses, outperforming all other methods. This consistent performance between ASGO and Muon aligns well with our discussions in Section 6. Moreover, The lightweight DASGO optimizer achieves competitive results against AdamW in both training and test loss metrics. DASGO demonstrates particularly strong performance during the initial training phase (first 200 training steps, which is about the first two epochs), but eventually exhibits a performance gap with ASGO. We discuss this performance gap in Section 7.1.
@@
To ensure optimal performance for each optimizer, we conducted a comprehensive HP search using random search [Bergstra and Bengio, 2012, Choi et al., 2019] within the following predefined ranges: initial lr: $10^{-5},10^{-1}]$; $\beta_{1}$: $[0.7,0.99]$; $\beta_{2}$: $[0.7,0.99]$; and selected the warmup factor from the set $[0.1,0.15,0.2,0.25,0.3]$. For the optimizers Shampoo and ASGO, we selected the update frequency $\tau$ from the set $[5,10,15]$ and for Shampoo, we selected its IO as either $\frac{1}{2}$ or $\frac{1}{4}$. HPs were selected based on validation loss after 20 epochs of training. For each optimizer, we performed a random search for 16 hours to find the optimal HP combination. These best HP choices for training NanoGPT are given in Table 5.
@@ C.2 Pretraining GPT2
The model configuration consists of 12 Transformer layers, 12 attention heads, and an embedding dimension of 768. The total number of training steps was 2400, with a batch size per GPU of 64. The model was trained using Distributed Data Parallel (DDP) on 4 NVIDIA GH200 GPUs, employing gradient accumulation over 8 steps. To ensure a fair comparison across all optimization algorithms, a consistent learning rate schedule was utilized: a linear warmup for the first 200 steps, followed by a cosine annealing decay to a final learning rate of $1\times 10^{-5}$ for the remainder of the training process. Training was conducted on the OpenWebText dataset with a sequence length of 512 tokens. For Shampoo optimizer, we used the Adam update grafting mentioned in [Shi et al., 2023, Eschenhagen et al., 2025]. For Muon, we used the Adam update RMS norm alignment in [Liu et al., 2025a]. We ran both the Newton-Schultz and the PolarExpress algorithms for 10 steps to compute the inverse square root of $V_{t}$ in the ASGO optimizer using the following parameters:
  • Newton-Schultz: (a,b,c)=(2,1.5,0.5)(a,b,c)=(2,-1.5,0.5)
  • PolarExpress:
    coeffs_list = [
    (8.28721201814563, -23.595886519098837, 17.300387312530933),
    (4.107059111542203, -2.9478499167379106, 0.5448431082926601),
    (3.9486908534822946 , -2.908902115962949, 0.5518191394370137),
    (3.3184196573706015 , -2.488488024314874, 0.51004894012372),
    (2.300652019954817, -1.6689039845747493, 0.4188073119525673),
    (1.891301407787398, -1.2679958271945868, 0.37680408948524835),
    (1.8750014808534479 , -1.2500016453999487, 0.3750001645474248),
    (1.875, -1.25, 0.375),
    (1.875, -1.25, 0.375),
    (1.875, -1.25, 0.375)]
@@
  • Learning rate (lr): $1\times 10^{-5},1\times 10^{-4},5\times 10^{-4},1\times 10^{-3},5\times 10^{-3},1\times 10^{-2},5\times 10^{-2},0.1]$
@@
Optimizer Learning Rate $\beta_{1}$ $\beta_{2}$ Weight Decay Damping $\epsilon$ Update Freq. ($\tau$)
Muon 0.01 0.9 N/A 0.1 N/A N/A
AdamW 0.005 0.9 0.99 0.1 $1\times 10^{-8}$ N/A
Shampoo 0.005 0.9 0.99 0.1 $1\times 10^{-10}$ 1
DASGO 0.01 0.9 0.9 0.1 $1\times 10^{-8}$ N/A
ASGO 0.01 0.9 0.8 0.1 $1\times 10^{-10}$ 1
(a) Train Loss
(b) Test Loss
Figure 6: Pretraining Performance on GPT2
@@ C.3 Finetune GPT2-Large
For the fine-tuning experiments on WikiText-2, our hyperparameter tuning focused exclusively on identifying the optimal learning rate for each optimizer under both the CLM and FIM objectives. We performed a grid search over a predefined set of learning rate candidates: $1\times 10^{-6},5\times 10^{-5},1\times 10^{-4},5\times 10^{-4},1\times 10^{-3},5\times 10^{-3}]$. The optimal learning rate for each optimizer and fine-tuning objective combination was determined by selecting the learning rate that yielded the lowest validation perplexity on the WikiText-2 validation set after the full 2 epochs of fine-tuning. Table 7 shows the optimal learning rate.
@@
CLM $5\times 10^{-5}$ $5\times 10^{-4}$ $\times 10^{-3} 1×1$$\times 10^{-3}$
FIM $1\times 10^{-4}$ $5\times 10^{-4}$ $\times 10^{-3} 1×1$$\times 10^{-3}$
@@
$\text{if }m<n\text{ then }V_{t}=\beta_{2}V_{t-1}+\left(1-\beta_{2}\right)M_{t}M_{t}^{\top}$ (4)
$\text{else }V_{t}=\beta_{2}V_{t-1}+\left(1-\beta_{2}\right)M_{t}^{\top}M_{t}$
$\text{if }m<n\text{ then }W_{t+1}=W_{t}-\eta_{t}(V_{t}+\epsilon I)^{-\frac{1}{2}}M_{t}$
$\text{else }W_{t+1}=W_{t}-\eta_{t}M_{t}(V_{t}+\epsilon I)^{-\frac{1}{2}}$
@@
  • Modified ASGO Update Direction (Coupled Newton, $\epsilon=10^{-8}$): $\Delta W^{\text{ASGO}}_{\text{CN},\epsilon=10^{-8}}=(V_{t}+10^{-8}I)^{-\frac{1}{2}}M_{t}$, where $(V_{t}+10^{-8}I)^{-\frac{1}{2}}$ is computed using the Coupled Newton algorithm for 50 iteration steps. Higham [2008a], Shi et al. [2023]
  • Modified ASGO Update Direction (Newton-Schulz, $\epsilon=10^{-8}$): $\Delta W^{\text{ASGO}}_{\text{NS},\epsilon=10^{-8}}=(V_{t}+10^{-8}I)^{-\frac{1}{2}}M_{t}$, where $(V_{t}+10^{-8}I)^{-\frac{1}{2}}$ is computed using the Newton-Schulz algorithm for 50 iteration steps. [^3] The Newton-Schulz algorithm is derived from Algorithm 6.35 on Page 153 in Higham [2008a]. We utilized the quintic version with parameters (a,b,c)=(2,1.5,0.5)(a,b,c)=(2,-1.5,0.5) as described in Jordan et al. [2024]. Higham [2008a], Bernstein and Newhouse [2024b], Jordan et al. [2024]
@@ Appendix E Proof of Nonsmooth Convergence of ASGO
$\text{Full-Matrix AdaGrad:}\quad\mathcal{O}\left(D_{\mathrm{F}}\sum_{j=1}^{m}\sum_{i=1}^{n}\sqrt{\sum_{t=0}^{T-1}[G_{t}]_{i,j}^{2}}\right)$
$\text{Shampoo:}\quad\mathcal{O}\left(\sqrt{r_{G}}D_{\mathrm{F}}\cdot{\rm tr}\left(\left(\sum_{t=0}^{T-1}G_{t}G_{t}^{\top}\right)^{\frac{1}{4}}\right)\cdot{\rm tr}\left(\left(\sum_{t=0}^{T-1}G_{t}^{\top}G_{t}\right)^{\frac{1}{4}}\right)\right)$
$\text{ASGO:}\quad\mathcal{O}\left(D_{\mathrm{op}}\cdot{\rm tr}\left(\left(\sum_{t=0}^{T-1}G_{t}G_{t}^{\top}\right)^{\frac{1}{2}}\right)\right)\leq\mathcal{O}\left(D_{\mathrm{op}}\cdot\sum_{j=1}^{m}\sqrt{\sum_{i=1}^{n}\sum_{t=0}^{T-1}[G_{t}]_{i,j}^{2}}\right).$
@@ Appendix G Proof of Section 6
Appendix H Incorporating Projection in ASGO
To achieve fully theoretically rigorous convergence bounds for ASGO, i.e., with $D_{\mathrm{op}}$ and $D_{\mathrm{F}}$ strict constants as $T$ increases, one well-established way is to incorporate a projection operation in the Algorithm [Hazan et al., 2007, Duchi et al., 2011]. Consider a closed convex set $\mathcal{W}\subseteq\mathbb{R}^{m\times n}$ with an optimum $W_{*}$ inside and diameters $\tilde{D}_{\mathrm{op}},\tilde{D}_{\mathrm{F}}$ such that for any $W_{1},W_{2}\in\mathcal{W}$, it holds that
$\left\|W_{1}-W_{2}\right\|_{\mathrm{op}}\leq\tilde{D}_{\mathrm{op}}\quad\text{and}\quad\left\|W_{1}-W_{2}\right\|_{\mathrm{F}}\leq\tilde{D}_{\mathrm{F}}.$
For Algorithm 1, we desire to search solution inside $\mathcal{W}$ and change the update step line 8 to
$Z_{t+1}=W_{t}-\eta_{t}\Lambda_{t}^{-1}G_{t}\quad\text{and}\quad W_{t+1}=\Pi_{\Lambda_{t}}^{\mathcal{W}}\left(Z_{t+1}\right),$
where the anisotropic projection operator $\Pi_{\Lambda_{t}}^{\mathcal{W}}:\mathbb{R}^{m\times n}\to\mathcal{W}$, as also incorporated in Hazan et al. [2007], Duchi et al. [2011], is defined as
$\Pi_{\Lambda_{t}}^{\mathcal{W}}\left(X\right)\triangleq\underset{Y\in\mathcal{W}}{\arg\min}\left\|X-Y\right\|_{\Lambda_{t}}^{2}$
for any $X\in\mathbb{R}^{m\times n}$, which enjoys the non-expansiveness of a projection operator:
$\left\|\Delta W_{t+1}-W_{*}\right\|_{\Lambda_{t}}^{2}={\rm tr}\left(\Delta W_{t+1}^{\top}\Lambda_{t}\Delta W_{t+1}\right)\leq{\rm tr}\left(\Delta Z_{t+1}^{\top}\Lambda_{t}\Delta Z_{t+1}\right)=\left\|\Delta Z_{t+1}\right\|_{\Lambda_{t}}^{2}$ (12)
with $\Delta W_{t+1}\triangleq W_{t+1}-W_{*}$ and $\Delta Z_{t+1}\triangleq Z_{t+1}-W_{*}$ since $W_{*}\in\mathcal{W}$. Substituting this non-expansiveness property in the proof of nonsmooth convergence, or more specifically, in (6), we can obtain that
$2\sum_{t=0}^{T-1}\mathbb{E}[f(W_{t})]-f(W_{*})$
$\leq$ $\frac{1}{\eta}\mathbb{E}\left[\sum_{t=0}^{T-1}{\rm tr}\left(\Delta W_{t}^{\top}\Lambda_{t}\Delta W_{t}-\Delta Z_{t+1}^{\top}\Lambda_{t}\Delta Z_{t+1}\right)\right]+\eta\mathbb{E}\left[\sum_{t=0}^{T-1}{\rm tr}\left(G_{t}^{\top}\Lambda_{t}^{-1}G_{t}\right)\right]$
[…]
$=$ $\frac{1}{\eta}\mathbb{E}\left[\sum_{t=0}^{T-1}{\rm tr}\left(\Delta W_{t}^{\top}\Lambda_{t}\Delta W_{t}-\Delta W_{t}^{\top}\Lambda_{t-1}\Delta W_{t}\right)\right]+\frac{1}{\eta}{\rm tr}\left(\Delta W_{0}^{\top}\Lambda_{-1}\Delta W_{0}\right)-\frac{1}{\eta}\mathbb{E}\left[{\rm tr}\left(\Delta W_{T}^{\top}\Lambda_{T-1}\Delta W_{T}\right)\right]$
$+\eta\mathbb{E}\left[\sum_{t=0}^{T-1}{\rm tr}\left(G_{t}^{\top}\Lambda_{t}^{-1}G_{t}\right)\right]$
$\leq$ $\frac{1}{\eta}{\rm tr}\left(\Delta W_{0}^{\top}\Lambda_{-1}\Delta W_{0}\right)+\frac{1}{\eta}\mathbb{E}\left[\sum_{t=0}^{T-1}{\rm tr}\left(\Delta W_{t}^{\top}(\Lambda_{t}-\Lambda_{t-1})\Delta W_{t}\right)\right]+\eta\mathbb{E}\left[\sum_{t=0}^{T-1}{\rm tr}\left(G_{t}^{\top}\Lambda_{t}^{-1}G_{t}\right)\right],$
which yields exactly the same result as (6). Since the rest of the proofs of both nonsmooth and smooth theory are based on (6), we can achieve the same convergence results as Theorems 1 and 3, while we have the upper bound on $\max_{0\leq t\leq T-1}\left\|W_{t}-W_{*}\right\|_{\mathrm{op}}\leq\tilde{D}_{\mathrm{op}}$ and $\max_{0\leq t\leq T-1}\left\|W_{t}-W_{*}\right\|_{\mathrm{F}}\leq\tilde{D}_{\mathrm{F}}$, where the diameters $\tilde{D}_{\mathrm{op}}$ and $\tilde{D}_{\mathrm{F}}$ depend on the geometry of the convex set $\mathcal{W}$ and remain constants when $T$ increases.